# A Large Scale Analysis of Gender Biases in Text-to-Image Generative Models

### v1

@@ 4 Gender Bias Analysis Experiments
For each prompt, e.g. the Activities prompt group has 1405 prompts, we calculate the ratio of male and female images. Let $\mathcal{I}^{p}=\{I_{1}^{p},I_{2}^{p},\ldots,I_{n}^{p}\}$, $n\leq 200$ be the set of gendered images for prompt $p$ after filtering and $\mathcal{G}:\mathcal{I}\rightarrow\{\textrm{female},\textrm{male}\}$ the mapping of images to unique genders according to InternVL2-8B. Then, we define the female-gendered images $F(p)$ and male-gendered images $M(p)$ as
$F(p)$ :=$:=\{I\in\mathcal{I}^{p}\ |\ \mathcal{G}(I)=\textrm{female}\}$ (1)
$M(p)$ :=$:=\{I\in\mathcal{I}^{p}\ |\ \mathcal{G}(I)=\textrm{male}\}$ (2)
@@
This allows us to estimate the distribution of female ratios across activities for a given model, i.e. we present the distribution of values of $\mathcal{R}_{f}$ as a histogram in Fig. 3. Overall, we find that the models generate similar gender ratios across all prompts (see Appendix D for more details). However, we also observe that models tend to generate more male-gendered images, as also observed by [35, 99].
@@ 4.1 Activities and Contexts
In Fig. 3, we find a large number of activities in which the set of images is male-dominated with $\mathcal{R}_{f}$ close to zero. We say a prompt or a prompt cluster is female-dominated (male-dominated) if $\mathcal{R}_{f}\geq 0.7$ ($\mathcal{R}_{f}\leq 0.3$), i.e. 70% (30%) or more (less) of images for this prompt or cluster are female-gendered. If $\mathcal{R}_{f}$ is between 50% and 70%, we speak of female-leaning clusters or prompts (equivalently male-leaning). While gender ratios of activities are distributed more evenly, the distribution of gender ratios for contexts is heavily skewed toward male images. This highlights a general trend to outputting males overall and men as the default. We calculate the top 10 (top 5) activities with the highest ratio of female- or male-gendered images across T2I models to showcase male- and female-dominated activities (contexts). We state the summary and the per-model average ratio of female- or male-gendered images of activities (contexts) in the given cluster for each activity (context) cluster. Results are in Fig. 4. For each cluster, we also report the average value of $\mathcal{R}_{f}$ across models (^$\hat{\mathcal{R}}_{f}$).
@@
Most female-dominated activities. Common female-dominated activities are crafting (^$\hat{\mathcal{R}}_{f}\approx 85\%$), which comprises activities such as “crocheting” or “making bracelets”, and pet-related activities (cat (^$\hat{\mathcal{R}}_{f}\approx 79\%$), pet (^$\hat{\mathcal{R}}_{f}\approx 81\%$)). birthday (^$\hat{\mathcal{R}}_{f}\approx 83\%$) contains activities related to parties. Further care-related activities are baby care (^$\hat{\mathcal{R}}_{f}\approx 88\%$) and volunteer (^$\hat{\mathcal{R}}_{f}\approx 82\%$), which are both examples of helping other people. The prominence on care-related activities reflects gender norms of women as caretakers and resembles human stereotypes, as women are described as “warm”, “sensitive to others”, and specifically “interested in children” [73]. The remaining highly female-dominated clusters in Fig. 4 are shopping-related activities (shopping (^$\hat{\mathcal{R}}_{f}\approx 83\%$), grocery (^$\hat{\mathcal{R}}_{f}\approx 76\%$)) and yoga-related activities in physical (^$\hat{\mathcal{R}}_{f}\approx 83\%$). Shopping-related activities include both shopping for daily necessities and clothes. Grocery shopping is known to be a household activity typically performed by women [22], and clothes are associated with women also in other prompt groups, especially contexts and objects (Sec. 4.2). Yoga was found to be seen as strongly female-typed [60]. Finally, “baking” (^$\hat{\mathcal{R}}_{f}\approx 76\%$) is a female-typed way of cooking [71].
@@
Most male-dominated activities. One common male-dominated activity type in Fig. 4 is outdoor household work (mowing (^$\hat{\mathcal{R}}_{f}\approx 3\%$), yard work (^$\hat{\mathcal{R}}_{f}\approx 16\%$)), which includes “mowing the lawn”, “cutting wood”, and “raking leaves”. Mowing the lawn specifically was identified as an activity typically performed by men [22]. Further male-dominated activities are car-related (car, car maintenance (both ^$\hat{\mathcal{R}}_{f}\approx 8\%$)), which is also male-typed [22], as well as media consumption, such as (computer) gaming (^$\hat{\mathcal{R}}_{f}\approx 9\%$) and movie watching (^$\hat{\mathcal{R}}_{f}\approx 10\%$) or tv viewing (^$\hat{\mathcal{R}}_{f}\approx 13\%$). According to [43], young men, on average, devote more time to “watching TV and video” and “computer games” than women of the same age. However, [82, 65] find that (computer) gamers being predominantly male is more of a stereotype than reality, meaning that T2I models perpetuate the marginalization of women in e-sports [65]. Smoking (^$\hat{\mathcal{R}}_{f}\approx 8\%$) explicitly refers to cannabis consumption, which, alongside other drug consumption including alcohol, is more common among men than women [91, 77, 39]. Football (^$\hat{\mathcal{R}}_{f}\approx 2\%$) is a strongly male-typed sport [68] and also strongly male-dominated in T2I models. Many male-gendered images in the parent-son (^$\hat{\mathcal{R}}_{f}\approx 2\%$) cluster are less surprising as prompts contain gendered words, i.e. “son”.
@@
Most female-dominated contexts. The only consistently female-leaning clusters are school (^$\hat{\mathcal{R}}_{f}\approx 68\%$) and student living (^$\hat{\mathcal{R}}_{f}\approx 55\%$), describing university environments, such as “classroom”, as well as childcare (^$\hat{\mathcal{R}}_{f}\approx 60\%$), containing places such as “playroom”. Teachers is a profession with female majority [64], which could explain the association of women with school places. Retail (^$\hat{\mathcal{R}}_{f}\approx 55\%$) contains various shopping locations, of which only a subset is strongly female-dominated. Such locations are related to fashion, such as “jewelry shop” or “hat shop”. Details on the retail cluster are in Sec. E.1.
@@
Most male-dominated contexts. In contrast to female-dominated contexts, the most male-dominated clusters focus on transportation (auto (^$\hat{\mathcal{R}}_{f}\approx 7\%$), shipping (^$\hat{\mathcal{R}}_{f}\approx 10\%$), parking and storage (^$\hat{\mathcal{R}}_{f}\approx 10\%$)) and industrial places (work (^$\hat{\mathcal{R}}_{f}\approx 9\%$)) Another strongly male-dominated cluster is baseball (^$\hat{\mathcal{R}}_{f}\approx 7\%$), which contains “baseball field” and various locations therein, such as “batting cage” or “pitcher’s mound”. We note that baseball is a male-leaning sport [60]. It is clear that T2I models do not associate women with industrial, work-related places, and places where women are depicted seem to be social places, such as schools or certain shops. In Sec. F.2, we provide further analyses of workplace gender bias.
@@ 4.2 Objects and Occupations
In Fig. 3, we observe a roughly Gaussian distribution for objects, peaking at a 0.4 female ratio. There are relatively few objects where models generate exclusively male- or female-gendered images. Gender distributions for occupations are highly polarized, with most occupations yielding only male-gendered images. Compared to actual work participation statistics, this reflects a clear bias amplification, in line with the findings in [80]. However, it also means that most previous work focusing on occupations has studied a particularly extreme example of bias amplification in T2I models. This further justifies our focus on activities and other everyday contexts. We list the top 5 object (occupation) clusters by female- and male ratios for each T2I model in Fig. 5.
@@
Most female-dominated objects. Main theme in female-dominated objects is clothing and accessories. Adorned personal (^$\hat{\mathcal{R}}_{f}\approx 77\%$) contains different types of jewelry, such as “crystal” or “ring”. Furniture (^$\hat{\mathcal{R}}_{f}\approx 64\%$) and textile (^$\hat{\mathcal{R}}_{f}\approx 62\%$) also fit this category, containing soft and textile-related objects such as “pillow” or “silk”. “Intimate wear” (^$\hat{\mathcal{R}}_{f}\approx 83\%$) contains underwear and swimwear. An additional female-leaning cluster, particularly in SD models, is “fruit” (^$\hat{\mathcal{R}}_{f}\approx 63\%$).
@@
Most male-biased objects. Common male-dominated objects are audio speakers (audio, ^$\hat{\mathcal{R}}_{f}\approx 28\%$) and music instruments (music, ^$\hat{\mathcal{R}}_{f}\approx 27\%$), vehicles (transportation, ^$\hat{\mathcal{R}}_{f}\approx 26\%$), and metal objects (metal, ^$\hat{\mathcal{R}}_{f}\approx 21\%$). We see a clear contrast between female-dominated objects, which are fashion-related, and male-dominated objects, which are technical. The male dominance in musical instruments is unexpected, as they oppose existing gendered associations of certain musical instruments [4, 3]. While we see these gender associations reflected in higher female ratios relative to other instruments (see Sec. E.2), musical instruments remain male-dominated.
As shown in Fig. 3, occupation prompts lead to extreme ratios of generated female- or male-gendered images. There are comparatively few female-dominated occupations, while most occupations are strongly male-dominated.
@@
Most female-dominated occupations. We find many relations to previously discussed female-dominated prompt clusters. For example, veterinary jobs (^$\hat{\mathcal{R}}_{f}\approx 90\%$) echo the pet-care-related activities, which we show are strongly female-dominated in Sec. 4.1. Nursing jobs (^$\hat{\mathcal{R}}_{f}\approx 96\%$), care jobs (^$\hat{\mathcal{R}}_{f}\approx 86\%$), and social jobs (^$\hat{\mathcal{R}}_{f}\approx 90\%$) all describe human-centered caring activities. Dental jobs (^$\hat{\mathcal{R}}_{f}\approx 84\%$) contains 4 occupations: “dental hygienist” and “dental assistant” are strongly female-dominated across all T2I models. “Dental technician” and “dentist” are female-dominated for all models except Flux-Schnell and SD-3.5-Large (see Sec. E.3 for detailed values). In the U.S., “dental hygienist” and “dental assistant” are occupations where $>90\%$ of the workforce are women, whereas $\approx 59\%$ of dental technicians and $\approx 40\%$ of dentists are women [64]. In Sec. F.3, we discuss bias amplification in T2I models regarding occupations by analyzing how gender ratios in images generated by models compare to actual workforce representation. Specifically, in Tab. 7, we show that T2I models not only replicate but often amplify existing gender skews found in occupations.
@@
Most male-dominated occupations. While many occupations are male-dominated, the most male-dominated occupations are blue-collar jobs involving physical labor. This is, for example, the case with installer jobs (^$\hat{\mathcal{R}}_{f}\approx 0\%$), construction jobs (^$\hat{\mathcal{R}}_{f}\approx 1\%$), and wood jobs (^$\hat{\mathcal{R}}_{f}\approx 1\%$). Wood jobs comprises occupations such as “carpenter” and “sawing machine operator”. Similarly, vehicle jobs (^$\hat{\mathcal{R}}_{f}\approx 1\%$) and rail jobs (^$\hat{\mathcal{R}}_{f}\approx 1\%$) are transportation industry occupations. In contrast to female-dominated occupations, none of the top male-dominated occupations are human-centric occupations.
@@ 4.3 Special Topics
We now look closely at specific topics within the activity prompt group that are particularly relevant to societal impact, namely household and work/money-related activities.
@@
Figure 6: (Top) Household-related clusters of activity prompts. (Bottom) Work and money-making related clusters of activity prompts.
@@
The division of household chores between spouses in heterosexual marriages is strongly moderated by gender [22, 44, 14, 51] and is relatively constant over time [26]. To select a subset of household-chores-related clusters, we classify all prompts in the activities prompt group as representing a household chore or not by an LLM (see Sec. F.1), specifically Phi-4. We cluster the resulting 105 activities and get 14 clusters, which we label manually and plot in Fig. 6 (Top).
@@
The seven clearly female-leaning clusters are laundry (^$\hat{\mathcal{R}}_{f}\approx 68\%$), dishes (^$\hat{\mathcal{R}}_{f}\approx 66\%$), making bed (^$\hat{\mathcal{R}}_{f}\approx 65\%$), clean bathroom (^$\hat{\mathcal{R}}_{f}\approx 62\%$), organizing (^$\hat{\mathcal{R}}_{f}\approx 57\%$), cleaning/vacuuming (^$\hat{\mathcal{R}}_{f}\approx 56\%$), and sweeping floor (^$\hat{\mathcal{R}}_{f}\approx 56\%$). Note that models are not uniformly biased in the categories, but generally, SD-3.5-Large and Flux-Schnell exhibit fewer biases than other T2I models in that there are more men in images representing these tasks. All these clusters are related to various forms of cleaning. If we compare with the typical household chore division [22], we find that most female-typed and shared cleaning chores are female-dominated in T2I models, e.g., “making bed” and “vacuuming” are listed as shared chores in [22], but “making bed” is female-dominated in images from by Flux variants. “Vacuuming” is female-dominated in SD variants.
@@
On the other end of the spectrum, we find the male-dominated clusters outside work (^$\hat{\mathcal{R}}_{f}\approx 5\%$), containing activities such as “working on the house”, and mowing lawn (^$\hat{\mathcal{R}}_{f}\approx 15\%$). Both are also more frequently performed by men [22]. This is also true for trash/cleaning (^$\hat{\mathcal{R}}_{f}\approx 42\%$), which, however, in our case, is a heterogenous cluster and also contains “doing a ton of spring cleaning” and “doing daily housework”, which are female-dominated in T2I models, while trash-related activities are strongly male-dominated. Interestingly, watering lawn (^$\hat{\mathcal{R}}_{f}\approx 46\%$) is male-typed in [22], but not clearly male-dominated in T2I models. Other not strongly gender-associated clusters are also listed as shared in [22].
Work and Money-Making. To assess gender bias regarding work or money-making-related activities, we also classify all 1405 activities by Phi-4 (see Sec. F.1) and cluster the resulting 139 prompts. This results in 20 clusters, which we label manually and show in Fig. 6 (Bottom).
No cluster other than teaching (^$\hat{\mathcal{R}}_{f}\approx 63\%$), work with animals (^$\hat{\mathcal{R}}_{f}\approx 62\%$), and writing (^$\hat{\mathcal{R}}_{f}\approx 59\%$) contains a majority of female-dominated activities. The cluster with the highest ratio of female-gendered images is teaching, which reflects our previous finding that teachers are associated with women. As already seen in Sec. 4.1, pet-related activities are frequently associated with women, and linking women with writing resembles the finding that humans associate women more than men with arts [63, 13]. Most other money-making activities, including regular work (^$\hat{\mathcal{R}}_{f}\approx 34\%$) and money-making (^$\hat{\mathcal{R}}_{f}\approx 24\%$), which refers to general activities related to money such as “worrying about money and time”, are male-dominated. We only see higher female ratios for job-seeking activities, i.e. job application (^$\hat{\mathcal{R}}_{f}\approx 46\%$) and job interview (^$\hat{\mathcal{R}}_{f}\approx 41\%$). This is concerning as underrepresenting women in work- and business-related contexts could reinforce existing stereotypes about women’s role in the workforce, perpetuating or even amplifying limiting gender norms of women as caretakers and men as breadwinners.
5 Discussion and Limitations
Analyzing many prompts and corresponding concepts, a clear picture emerges of which activities, contexts, objects, and occupations T2I models associate most with women or men. First of all, T2I models tend to generate more images of men than women. This highlights how men are the general default in model outputs, including in gender-neutral prompts and categories. The scenarios where men are not the default are in prompts of activities, contexts, objects, and occupations that are stereotypically feminine. We consistently observe that scenarios with a high rate of female-gendered images portray women in traditional roles: as homemakers, while shopping, or engaged in arts and beauty in our activity prompts; as caring and service-oriented in our contexts and occupations; and with fashionable and soft objects. This contrasts with male-dominated scenarios. Men are associated with physical work, both in the household and at their jobs, working with machinery, and are strongly associated with business. While this reflects the greater numbers of women in caretaking roles and men in machinery-related or business leadership roles that exist in society, our analysis clearly shows that gender stereotypes are replicated, reinforced, and further amplified in T2I models. While we can only numerically prove bias amplification in occupations due to the existence of workforce labor statistics to compare against which do not exist for other scenarios (see Sec. F.3), we hypothesize that the same is happening for other scenarios beyond occupations.
Our study offers detailed and valuable insights into gender bias in T2I models, which should be revisited and highlighted repeatedly [2]. However, our focus on binary gender leaves room for further exploration of social biases in T2I models. Regarding validity, we take concrete steps to include only prompts and images suitable for evaluating gender bias through person detection and filtering. Although our analysis relies on automatic methods, the strength of the evidence supporting our conclusion leaves little doubt that the observed effects stem from spurious biases in automatic processing. This study does not examine how prompts that evoke particular gender identities (e.g., a female engineer) can solve some of these issues, or whether such outputs have other gendered stereotypes. Rather, our contribution is analyzing outputs for gender-neutral prompts to unpack underlying defaults and gender biases present in models.

### v2

@@ 4 Gender Bias Analysis Experiments
For each prompt, e.g. the Activities prompt group has 1405 prompts, we calculate the ratio of male and female images. Let $\mathcal{I}^{p}=\{I_{1}^{p},I_{2}^{p},\ldots,I_{n}^{p}\}$, $n\leq 200$ be the set of gendered images for prompt $p$ after filtering and $\mathcal{G}:\mathcal{I}\rightarrow\{\textrm{female},\textrm{male}\}$ the mapping of images to unique genders according to InternVL2-8B. Then, we define the female- and male-gendered images $F(p)$ and $M(p)$ as
$F(p)$ $:=\{I\in\mathcal{I}^{p}\ |\ \mathcal{G}(I)=\textrm{female}\}$ (1)
$M(p)$ $:=\{I\in\mathcal{I}^{p}\ |\ \mathcal{G}(I)=\textrm{male}\}$ (2)
@@
This allows us to estimate the distribution of female ratios across activities for a given model, i.e. we present the distribution of values of $\mathcal{R}_{f}$ as a histogram in Fig. 2. Overall, we find that the models generate similar gender ratios across all prompts (see Section F.1 for more details). However, we also observe that models tend to generate more male-gendered images, as also observed in (Ghosh & Caliskan, 2023; Zhang et al., 2023).
@@ 4.1 Activities and Contexts
In Fig. 2, we find a large number of activities in which the set of images is male-dominated with $\mathcal{R}_{f}$ close to zero. We say a prompt or a prompt cluster is female-dominated (male-dominated) if $\mathcal{R}_{f}\geq 0.7$ ($\mathcal{R}_{f}\leq 0.3$), i.e. 70% (30%) or more (less) of images for this prompt or cluster are female-gendered. If $\mathcal{R}_{f}$ is between 50% and 70%, we speak of female-leaning clusters or prompts (equivalently male-leaning). While gender ratios of activities are distributed more evenly, the distribution of gender ratios for contexts is heavily skewed toward male images. This highlights a general trend to outputting males overall and men as the default. We calculate the top 10 (top 5) activities with the highest ratio of female- or male-gendered images across T2I models to showcase male- and female-dominated activities (contexts). We state the summary and the per-model average ratio of female- or male-gendered images of activities (contexts) in the cluster for each activity (context) cluster. Results are in Fig. 3. For each cluster, we also report the mean $\mathcal{R}_{f}$ across models ($\hat{\mathcal{R}}_{f}$).
Figure 3: (Top) Top 10 most female-dominated (top row) and top 10 most male-dominated (bottom row) activity clusters. Bars indicate ratio of female-gendered images generated from 5 T2I models averaged over prompts in each cluster. Error line indicates the std. dev. across prompts. (Bottom) Top 5 most female-dominated (left) and top 5 most male-dominated (right) context clusters.
@@
Most female-dominated activities. Common female-dominated activities are crafting ($\hat{\mathcal{R}}_{f}\approx 85\%$), which comprises activities such as “crocheting” or “making bracelets”, and pet-related activities (cat ($\hat{\mathcal{R}}_{f}\approx 79\%$), pet ($\hat{\mathcal{R}}_{f}\approx 81\%$)). birthday ($\hat{\mathcal{R}}_{f}\approx 83\%$) contains activities related to parties. Further care-related activities are baby care ($\hat{\mathcal{R}}_{f}\approx 88\%$) and volunteer ($\hat{\mathcal{R}}_{f}\approx 82\%$), which are both examples of helping other people. The prominence on care-related activities reflects gender norms of women as caretakers and resembles human stereotypes, as women are described as “warm”, “sensitive to others”, and specifically “interested in children” (Rudman et al., 2012). The remaining highly female-dominated clusters in Fig. 3 are shopping-related activities (shopping ($\hat{\mathcal{R}}_{f}\approx 83\%$), grocery ($\hat{\mathcal{R}}_{f}\approx 76\%$)) and yoga-related activities in physical ($\hat{\mathcal{R}}_{f}\approx 83\%$). Shopping-related activities include both shopping for daily necessities and clothes. Grocery shopping is known to be a household activity typically performed by women (Coltrane, 1989), and clothes are also associated with women in other prompt groups, especially contexts and objects (Section 4.2). Yoga was found to be seen as strongly female-typed (Matteo, 1986). Finally, “baking” ($\hat{\mathcal{R}}_{f}\approx 76\%$) is a female-typed way of cooking (Rokicki et al., 2016).
@@
Most male-dominated activities. One common male-dominated activity type in Fig. 3 is outdoor household work (mowing ($\hat{\mathcal{R}}_{f}\approx 3\%$), yard work ($\hat{\mathcal{R}}_{f}\approx 16\%$)), which includes “mowing the lawn”, “cutting wood”, and “raking leaves”. Mowing the lawn specifically was identified as an activity typically performed by men (Coltrane, 1989). Further male-dominated activities are car-related (car, car maintenance (both $\hat{\mathcal{R}}_{f}\approx 8\%$)), which is also male-typed (Coltrane, 1989), as well as media consumption, such as (computer) gaming ($\hat{\mathcal{R}}_{f}\approx 9\%$) and movie watching ($\hat{\mathcal{R}}_{f}\approx 10\%$) or tv viewing ($\hat{\mathcal{R}}_{f}\approx 13\%$). According to Hilbrecht et al. (2008), young men, on average, devote more time to “watching TV and video” and “computer games” than women of the same age. However, Shaw (2012); Paaßen et al. (2017) find that (computer) gamers being predominantly male is more of a stereotype than reality, meaning that T2I models perpetuate the marginalization of women in e-sports (Paaßen et al., 2017). Smoking ($\hat{\mathcal{R}}_{f}\approx 8\%$) explicitly refers to cannabis consumption, which, alongside other drug consumption including alcohol, is more common among men than women (Wilsnack et al., 2000; Schulte et al., 2009; Hemsing & Greaves, 2020). Football ($\hat{\mathcal{R}}_{f}\approx 2\%$) is a strongly male-typed sport (Plaza et al., 2017) and also strongly male-dominated in T2I models. Many male-gendered images in the parent-son ($\hat{\mathcal{R}}_{f}\approx 2\%$) cluster are less surprising as prompts contain gendered words, i.e. “son”.
@@
Most female-dominated contexts. The only consistently female-leaning clusters are school ($\hat{\mathcal{R}}_{f}\approx 68\%$) and student living ($\hat{\mathcal{R}}_{f}\approx 55\%$), describing university environments, such as “classroom”, as well as childcare ($\hat{\mathcal{R}}_{f}\approx 60\%$), containing places such as “playroom”. Teachers is a profession with female majority (in the USA, see (Bureau of Labour Statistics, 2023)), which could explain the association of women with school places. Retail ($\hat{\mathcal{R}}_{f}\approx 55\%$) contains various shopping locations, of which only a subset is strongly female-dominated. Such locations are related to fashion, such as “jewelry shop” or “hat shop”. Details on the retail cluster are in Section H.1.
@@
Most male-dominated contexts. In contrast to female-dominated contexts, the most male-dominated clusters focus on transportation (auto ($\hat{\mathcal{R}}_{f}\approx 7\%$), shipping ($\hat{\mathcal{R}}_{f}\approx 10\%$), parking and storage ($\hat{\mathcal{R}}_{f}\approx 10\%$)) and industrial places (work ($\hat{\mathcal{R}}_{f}\approx 9\%$)). Another strongly male-dominated cluster is baseball ($\hat{\mathcal{R}}_{f}\approx 7\%$), which contains “baseball field” and various locations therein, such as “batting cage” or “pitcher’s mound”. We note that baseball is a male-leaning sport (Matteo, 1986). It is clear that T2I models do not associate women with industrial, work-related places, and places where women are depicted seem to be social places, such as schools or certain shops. In Section I.3, we provide further analyses of workplace gender bias.
@@ 4.2 Objects and Occupations
In Fig. 2, we observe a roughly Gaussian distribution for objects, peaking at a 0.4 female ratio. There are relatively few objects where models generate exclusively male- or female-gendered images. Gender distributions for occupations are highly polarized, with most occupations yielding only male-gendered images. Compared to actual work participation statistics, this reflects a clear bias amplification, in line with the findings by Seshadri et al. (2024). However, it also means that most previous work focusing on occupations has studied a particularly extreme example of bias amplification in T2I models. This further justifies our focus on activities and other everyday contexts. We list the top 5 object (occupation) clusters by female- and male ratios for each T2I model in Fig. 4.
@@
Most female-dominated objects. Main theme in female-dominated objects is clothing and accessories. Adorned personal ($\hat{\mathcal{R}}_{f}\approx 77\%$) contains different types of jewelry, such as “crystal” or “ring”. Furniture ($\hat{\mathcal{R}}_{f}\approx 64\%$) and textile ($\hat{\mathcal{R}}_{f}\approx 62\%$) also fit this category, containing soft and textile-related objects such as “pillow” or “silk”. “Intimate wear” ($\hat{\mathcal{R}}_{f}\approx 83\%$) contains underwear and swimwear. Another female-leaning cluster, particularly in SD models, is “fruit” ($\hat{\mathcal{R}}_{f}\approx 63\%$).
@@
Most male-biased objects. Common male-dominated objects are audio speakers (audio, $\hat{\mathcal{R}}_{f}\approx 28\%$) and music instruments (music, $\hat{\mathcal{R}}_{f}\approx 27\%$), vehicles (transportation, $\hat{\mathcal{R}}_{f}\approx 26\%$), and metal objects (metal, $\hat{\mathcal{R}}_{f}\approx 21\%$). We see a clear contrast between female-dominated objects, which are fashion-related, and male-dominated objects, which are technical. The male dominance in musical instruments is unexpected, as they oppose existing gendered associations of certain musical instruments (Abeles & Porter, 1978; Abeles, 2009). While we see these gender associations reflected in higher female ratios relative to other instruments (see Section H.2), musical instruments remain male-dominated.
@@
Most female-dominated occupations. We find many relations to previously discussed female-dominated prompt clusters. For example, veterinary jobs ($\hat{\mathcal{R}}_{f}\approx 90\%$) echo the pet-care-related activities, which we show are strongly female-dominated in Section 4.1. Nursing jobs ($\hat{\mathcal{R}}_{f}\approx 96\%$), care jobs ($\hat{\mathcal{R}}_{f}\approx 86\%$), and social jobs ($\hat{\mathcal{R}}_{f}\approx 90\%$) all describe human-centered caring activities. Dental jobs ($\hat{\mathcal{R}}_{f}\approx 84\%$) contains 4 occupations: “dental hygienist” and “dental assistant” are strongly female-dominated across all T2I models. “Dental technician” and “dentist” are female-dominated for all models except Flux-Schnell and SD-3.5-Large (see Section H.3 for detailed values). In the U.S., “dental hygienist” and “dental assistant” are occupations where $90\%$ of the workforce are women, whereas $\approx 59\%$ of dental technicians and $\approx 40\%$ of dentists are women.
@@
Most male-dominated occupations. While many occupations are male-dominated, the most male-dominated occupations are blue-collar jobs involving physical labor. This is, for example, the case with installer jobs ($\hat{\mathcal{R}}_{f}\approx 0\%$), construction jobs ($\hat{\mathcal{R}}_{f}\approx 1\%$), and wood jobs ($\hat{\mathcal{R}}_{f}\approx 1\%$). Wood jobs comprises occupations such as “carpenter” and “sawing machine operator”. Similarly, vehicle jobs ($\hat{\mathcal{R}}_{f}\approx 1\%$) and rail jobs ($\hat{\mathcal{R}}_{f}\approx 1\%$) are transportation industry occupations. In contrast to female-dominated occupations, none of the top male-dominated occupations are human-centric.
@@ 4.3 Special Topics
We now look closely at specific topics of the activity prompt group, which have particular societal impact, namely household activities and bias amplification. More analyses on work/money-related activities are in Section I.1 and on bias amplification in jobs in Section I.5.
@@
Figure 5: Household-related clusters of activity prompts.
@@
The division of household chores between spouses in heterosexual marriages is strongly moderated by gender (Coltrane, 1989; Hiller & Philliber, 1986; Cerrato & Cifre, 2018; Kroska, 2003) and is relatively constant over time (Douthitt, 1989). To select a subset of household-chores-related clusters, we classify all prompts in the activities prompt group as representing a household chore or not by an LLM (see Section I.2), specifically Phi-4. We cluster the resulting 105 activities and get 14 clusters, which we label manually and plot in Fig. 5.
@@
The seven clearly female-leaning clusters are laundry ($\hat{\mathcal{R}}_{f}\approx 68\%$), dishes ($\hat{\mathcal{R}}_{f}\approx 66\%$), making bed ($\hat{\mathcal{R}}_{f}\approx 65\%$), clean bathroom ($\hat{\mathcal{R}}_{f}\approx 62\%$), organizing ($\hat{\mathcal{R}}_{f}\approx 57\%$), cleaning/vacuuming ($\hat{\mathcal{R}}_{f}\approx 56\%$), and sweeping floor ($\hat{\mathcal{R}}_{f}\approx 56\%$). Note that models are not uniformly biased in the categories, but generally, SD-3.5-Large and Flux-Schnell exhibit fewer biases than other T2I models in that there are more men in images representing these tasks. All these clusters are related to various forms of cleaning. If we compare with the typical household chore division (Coltrane, 1989), we find that most female-typed and shared cleaning chores are female-dominated in T2I models, e.g., “making bed” and “vacuuming” are shared chores in (Coltrane, 1989), but “making bed” is female-dominated in images from by Flux variants. “Vacuuming” is female-dominated in SD variants.
@@
On the other end of the spectrum, we find the male-dominated clusters outside work ($\hat{\mathcal{R}}_{f}\approx 5\%$), containing activities, e.g. “working on the house”, and mowing lawn ($\hat{\mathcal{R}}_{f}\approx 15\%$). Both are more frequently performed by men (Coltrane, 1989). This is also true for trash/cleaning ($\hat{\mathcal{R}}_{f}\approx 42\%$) that in our case is a heterogenous cluster and also contains “doing a ton of spring cleaning” and “doing daily housework” that are female-dominated in T2I models, while trash-related activities are strongly male-dominated. Interestingly, watering lawn ($\hat{\mathcal{R}}_{f}\approx 46\%$) is male-typed in (Coltrane, 1989), but not clearly male-dominated in T2I models. Other not strongly gender-associated clusters are listed as shared in (Coltrane, 1989).
Bias Amplification in Activities. While previous work (Seshadri et al., 2024; Luccioni et al., 2024) has investigated bias amplification in occupations (we confirm these findings in Section I.5), we also show bias amplification of T2I models in activities. To study bias amplification in activities, we retrieve matching images for activity prompts from LAION-400M (Schuhmann et al., 2021) and examine the gender that is represented in this dataset.
Male majority Female majority
reduced amplified reduced amplified
Flux 12.68% 87.32% 60.49% 39.51%
Flux-Schnell 18.31% 81.69% 71.60% 28.40%
SD-3.5-Large 25.35% 74.65% 60.49% 39.51%
SD-3.5-Medium 35.21% 64.79% 40.74% 59.26%
SD-3-Medium 16.90% 83.10% 60.49% 39.51%
Table 3: Bias amplification for male-majority and female-majority activities wrt. LAION-400m.
We chose LAION-400m because it is representative of the web-scale datasets typically used to train T2I image models. We use a text-based method to find all images for which the non-stopword lemmas (extracted via spaCy) of the activity prompt are a subset of those from the image caption. If over 10,000 images match, we sample 10,000 randomly. We detect people using YOLOv10 and infer perceived gender with InternVL2-8B, following the setup in Sections 3.2 and E.1. Images without recognizable gender or with mixed genders are discarded, leaving 152 prompts with $\geq$ 50 matched images each. The average female ratio across these is  52%, while in T2I-generated images it’s  41%, showing underrepresentation of women in generations relative to LAION-400m. Given that LAION-400M is representative of data used to train T2I models, this is interesting: it suggests that it may not only be the training data that is leading to greater representation of men in outputs, but there is amplification of male representation in the model itself. This is significant, as much research on bias focuses on the training data.
We label activities as “female majority” or “male majority” based on LAION-400m proportions (for example: if cooking has more female representation in images in the dataset, it would be labeled “female majority”). We then assess whether the majority gender ratio increases (bias amplification) or decreases (bias reduction) in T2I outputs. For example, if there is greater female representation in images of cooking in T2I models than in the LAION-400m dataset, this would be labeled as “bias amplification”. As shown in Table 3, male-majority activities show increased male ratios, while female-majority ones show mixed outcomes but generally reduced female ratios. This indicates T2I models amplify male-gender bias beyond what is in training data, motivating deeper analysis of web-scale datasets. Further details on bias amplification in activities are in Section I.4.